# Ondřej Mokrý Dept. of Telecommunications

### v2

@@ Abstract
The paper focuses on inpainting missing parts of an audio signal spectrogram. The autoregression-based Janssen algorithm, the state-of-the-art for the time-domain audio inpainting, is adapted for the time-frequency setting. This novel method, termed Janssen-TF, is compared to the deep-prior neural network approach using both objective metrics and a subjective listening test, proving Janssen-TF to be superior in all the considered measures.
@@
audio inpainting, autoregression, deep prior, time-frequency, spectrogram.
@@ I Introduction
Audio inpainting, sometimes referred to as missing audio interpolation or concealment, is the task of filling in the absent parts of audio with a meaningful content. Typically, the missing information is in the form of lost samples in the time domain. If compact regions of such samples are present, they are often referred to as the gaps. For the described task, plenty of methods have been developed: those based on the autoregressive character of audio [1, 2], methods based on the sparsity of the time-frequency representation of audio [3, 4, 5, 6], or methods relying on similarities in audio recordings [7, 8, 9]. Quite recently, deep learning audio inpainting methods appeared [10, 11]. For gaps of up to 50 ms in length, the Janssen method [2, 12] has been evaluated the best for a long time, but PHAIN [13] may become a new state-of-the art.
The time-domain inpainting represents the most natural case from the point of view of the application to signal recovery. Nonetheless, many practical algorithms, such as classic audio codecs (MP3, AAC, etc.), operate on short-time signal spectra. From this perspective, it is valuable to think about inpainting in the time-frequency domain. [^1] In other fields, one can go even further and require inpainting purely in the frequency domain, see for example [14], [15]. In such a case, a portion of time-frequency (TF) coefficients222 typically in the form of a complex-valued matrix are missing, which need to be recovered to allow a later reconstruction of the time-domain signal. Indeed, there have been a few methods developed in the TF context, all of them relying on neural modeling [16, 17, 18]. The most recent work [18] is based on the concept of a “deep prior”, a neural network that does not need training in the most common sense.
Note that a compact gap in time bears some fundamental differences compared with a gap in the form of a missing number of consecutive columns of a spectrogram. This follows from the nature of time-frequency transforms, which use windows with overlaps to compute spectrograms, see for instance [19]. In particular, a gap in a spectrogram does not fully correspond to a time-domain gap and vice versa; there is always an amplitude fading in one of the domains thanks to the windowing.
@@
II Deep Prior Audio Inpainting (DPAI)
The deep prior, introduced in [20], is a concept used for the reconstruction of corrupted signals using a convolutional neural network (CNN). In contrast to most CNN applications, [20] does not utilize a training dataset; the reconstruction only involves a proper CNN architecture and the corrupted signal. The deep prior concept has also proven that the network architecture matters much more than previously thought.
II-A Application of Deep Prior to Audio Inpainting
@@
An encoder–decoder CNN denoted $f_{\theta}(\mathbf{n})$ is assumed to be able to generate any signal of size $N$ using a random but fixed noise input $\mathbf{n}\in\mathbb{R}^{N}$, using a proper $\theta$. This is a reliable assumption since the number of network parameters ${\theta}$ (weights and biases) is significantly greater than $N$.
In the deep prior reconstruction, parameters $\theta$ are initialized randomly. To obtain a reconstruction result, $\theta$ is then optimized as follows:
@@
where $E$ is a differentiable data fidelity function measuring the closeness of its two inputs. The optimal $\theta^{*}$ can be obtained using an iterative optimizer, e.g., Adam [21]. The crucial observation is that the optimization must be interrupted before it reaches $\theta^{*}$, for which we would have $f_{\theta^{*}\!}(\mathbf{n})=\mathbf{x}^{\textup{cor}}$, meaning basically overfitting. Instead, early stopping of the training is desired, resulting in an estimate $\hat{\theta}$ and a corresponding $\hat{\mathbf{x}}=f_{\hat{\theta}}(\mathbf{n})$ [20].
@@
The authors of [18] used an analog approach to inpaint audio in the transformed domain. Through the use of the short-time Fourier transform (STFT), denoted $\mathbf{L}$ in the following, an audio recording $\mathbf{x}$ is considered merely in the TF domain, where it forms a spectrogram matrix $\mathbf{L}\mathbf{x}=\mathbf{X}\in\mathbb{C}^{T\times F}$. Therefore the CNN operates on spectrograms and $\hat{\mathbf{X}}=f_{\hat{\theta}}(\mathbf{N})$ for a random but fixed matrix $\mathbf{N}$. Once $\hat{\mathbf{X}}$ has been established, the time-domain solution is obtained by the inverse STFT.
@@
Regarding the mask, in [18] only the case of completely missing columns of the spectrogram is considered, i.e., for the TF mask $\mathbf{M}\in\{0,1\}^{T\times F}$ it holds that each column is either full of ones or zeros. Nevertheless, a generalization to other TF masks is straightforward. The observed spectrogram $\mathbf{X}^{\textup{cor}}$ is finally modeled as $\mathbf{X}^{\textup{cor}}\!=\mathbf{M}\odot\mathbf{X}^{\textup{orig}}$, where $\mathbf{X}^{\textup{orig}}$ corresponds to $\mathbf{x}^{\textup{orig}}$ in the time domain, i.e., $\mathbf{X}^{\textup{orig}}=\mathbf{L}\mathbf{x}^{\textup{orig}}$. Note that by analogy with the time-domain case, we have the feasible set $\mathit{\Gamma}_{\textup{TF}}=\{\mathbf{X}\,|\,\mathbf{M}\odot\mathbf{X}=\mathbf{M}\odot\mathbf{X}^{\textup{orig}}\}$.
The loss function $E$ in (1), which is optimized with respect to $\theta$, is in the form
$E=\alpha_{1}\mathrm{MSE}(\mathbf{M}\odot\mathbf{X},\mathbf{X}^{\textup{cor}})+\alpha_{2}\mathrm{MSS}(\mathbf{M}\odot\mathbf{X},\mathbf{X}^{\textup{cor}}),$ (2)
where MSE is the Mean Squared Error and $\mathrm{MSS}$ is the multi-scale spectrogram loss [22]. In [18], the weights $\alpha_{1}$ and $\alpha_{2}$ have values 1 and 0.1 respectively.
II-B Architecture
@@
The U-Net is a well-known CNN architecture consisting of several convolutional layers, downsampling (encoder) and upsampling (decoder) operations. In certain applications including audio inpainting, the existence of skip connections between the encoder and decoder stages is important. The authors of [18] used an architecture similar to U-Net called MultiResUNet [23] with an additional inclusion of the so-called harmonic convolution [24], which proved to be the best among multiple tested architectures in [18]. The harmonic convolution is a variation on the common 2D convolution that takes into account higher harmonics of audio.
II-C Variants
@@
The DPAI reconstructs the whole $\hat{\mathbf{X}}$. Yet, considering that only a part of the spectrogram was unavailable, the authors of [18] evaluate the performance of their method after injecting the reliable coefficients back to the solution as a postprocessing step. Formally, such an action is a projection
@@
onto the set $\mathit{\Gamma}_{\textup{TF}}$. Note that a resulting sharp transition between the regions is smeared in the time domain due to the nature of the STFT [19]. We call the reconstructions modified by (3) with context, while the others using the entire $\hat{\mathbf{X}}$ are without context. In the experimental part, we will study whether such a modification brings a significant performance improvement.
III Janssen spectrogram inpainting (Janssen-TF)
@@
$\sum_{i=1}^{p+1}a_{i}x_{n+1-i}=e_{n},\quad n=1,\dots,N+p,$ (4)
where $\mathbf{a}=[1,a_{2},\dots,a_{p+1}]^{\top}$ is the vector of model parameters (AR coefficients), $p$ is the model order, and $\mathbf{x}$ is zero-padded such that its final length is $N+p$. The error term $\mathbf{e}\in\mathbb{R}^{N+p}$ represents a realization of a zero-mean white noise process with variance $\sigma^{2}$ [25, Def. 3.1.2].
@@
Step 1 can be solved efficiently using standard methods, such as the Durbin–Levinson recursion (see e.g. the function lpc in Matlab). Regarding step 2, we propose to adapt the Janssen iteration to the TF setting, denoted Janssen-TF. In our adaptation, the signal spectrogram is constrained to fit the observation, $\mathbf{X}^{\textup{cor}}$, while the signal itself follows the estimated AR model. To formalize this modification, note that the error in (4) can be written as $\mathbf{e}=\mathbf{A}\mathbf{x}$, where $\mathbf{A}\in\mathbb{R}^{(N+p)\times N}$ is a Toeplitz matrix composed using the (fixed) coefficients $\mathbf{a}$. We can then pose the proposed modification of step 2 as
@@
where $\mathbf{L}$ is the STFT and $\iota_{\mathit{\Gamma}_{\textup{TF}}}$ denotes the indicator function of the convex set $\mathit{\Gamma}_{\textup{TF}}$ of feasible spectrograms. [^3] Note that the original Janssen algorithm minimizes the function $\frac{1}{2}\|\mathbf{A}\mathbf{x}\|+\iota_{\mathit{\Gamma}_{\textup{T}}}(\mathbf{x})$ in subproblem 2. As such, the problem (5) is convex and it is solvable using standard numerical methods. We propose to employ the alternating direction method of multipliers (ADMM), which includes two principal steps [26, Sec. 3.1.1]:
@@
The whole algorithm is summarized in Alg. 1. For the simplification of (6a), we have posed the assumption [^4] This assumption is met for instance if $\mathbf{L}$ corresponds to a Parseval tight frame, which can be ensured by a suitable setting of the STFT. $\mathbf{L}^{*}\mathbf{L}=\mathbf{I}$, leading to $\bar{\mathbf{x}}^{(k+1)}=\operatorname{prox}_{\frac{1}{2}\|\cdot\|^{2}}(\mathbf{L}^{*}(\mathbf{z}^{(k)}-\mathbf{u}^{(k)}))$, see [27, Rem. 3], which then corresponds to line 1 of Alg. 1 [28, Table 1]. Similarly, (6b) boils down to line 1 of the algorithm; this is computed via (3). The hyperparameters of the algorithm are the number of outer and inner iterations ($I$ and $K$, respectively), and the step size $\rho>0$.
@@ 11      $\mathbf{x}^{(i+1)}=\bar{\mathbf{x}}^{(K+1)}$
Note that several algorithms can be used to solve (5) numerically, such as the primal–dual algorithm [29] or the Douglas–Rachford or the forward–backward algorithms [28], combined with the approximal operator of the function $\iota_{\mathit{\Gamma}_{\textup{TF}}}(\mathbf{L}\,\cdot)$ [30]. However, the variant with ADMM described above, denoted Janssen-TF-ADMM, exhibited the best performance out of six algorithms in our preliminary testing.
IV Experiments
IV-A Data, setup and metrics
@@
The STFT (i.e., the operator $\mathbf{L}$) utilizes a 2048-sample-long Hann window (which corresponds to 128 ms), with 75% overlap and 2048 frequency channels; this setting is the default for the time-frequency transform of the librosa package [^6] https://librosa.org/doc/0.10.2/generated/librosa.stft.html, as used in [18], and ensures that $\mathbf{L}$ forms a Parseval tight frame.
@@
  • gapwise Janssen method in the time domain [12].
The gapwise method is a twist to the original Janssen algorithm [2], which, instead of processing by means of overlapping frames, makes use of a single segment consisting of the whole gap and its nearby context.
@@
IV-B Results
The objective comparison, including the proposed Janssen-TF method, is shown in Fig. 1 for the case of the DPAI dataset. The first observation is that the time-domain-based Janssen method [12] is not competitive with the other approaches. This can be explained by the incompatibility of gaps in the time and time-frequency domains (a sharp mask in the spectrogram domain affects a longer signal portion when transformed into the time domain, and the whole affected segment is treated as missing). While the average performance of the DPAI variants mostly lies within the confidence intervals of each other, suggesting limited significance of the differences, the proposed Janssen-TF-ADMM method largely outperforms the competitors in both SNR and ODG (except the longest gaps).
Figure 1: Comparison of the inpainting methods using objective metrics – SNR (top) and ODG (bottom). The plots show results averaged over the 8 test signals together with the bootstrap interval estimates of the mean values at the $\alpha=5\%$ significance level [35, p. 160]. If the intervals do not overlap, it may be concluded that the difference of the means is statistically significant.
To support the objective evaluation on the first dataset, Fig. 2 presents the results of the subjective listening test. Its results correlate well with the objective evaluation; in particular, Janssen-TF-ADMM is superior to the other methods. Moreover, this result is statistically significant in terms of medians. On the other hand, the DPAI variants scored almost the same.
anchorreferenceDPAI without contextDPAI with contextgapwise JanssenJanssen--TF-ADMM0020202020404040406060606080808080100100100100MUSHRA score

### v3

@@ Abstract
The paper focuses on inpainting missing parts of an audio signal spectrogram, i.e., estimating the lacking time-frequency coefficients. The autoregression-based Janssen algorithm, a state-of-the-art for the time-domain audio inpainting, is adapted for the time-frequency setting. This novel method, termed Janssen-TF, is compared with the deep-prior neural network approach using both objective metrics and a subjective listening test, proving Janssen-TF to be superior in all the considered measures.
@@
audio inpainting, autoregression, deep prior, DPAI, time-frequency, spectrogram.
@@ I Introduction
Audio inpainting, sometimes referred to as missing audio interpolation or concealment, is the task of filling in the absent parts of audio with a meaningful content. Typically, the missing information is in the form of lost samples in the time domain. If compact regions of such samples are present, they are often referred to as the gaps. For the described task, numerous methods have been developed: those exploiting the autoregressive character of audio [1, 2], methods based on the sparsity of the time-frequency representation of audio [3, 4, 5, 6], or methods relying on similarities in audio recordings [7, 8, 9]. Quite recently, deep learning audio inpainting methods appeared [10, 11]. For gaps of up to 50 ms in length, the Janssen method [2, 12] has been evaluated the best for a long time, but PHAIN [13] may become a new state-of-the art.
The time-domain inpainting represents the most natural case from the point of view of the application to signal recovery. Nonetheless, many practical algorithms, such as classic audio codecs (MP3, AAC, etc.), operate on short-time signal spectra. From this perspective, it is valuable to think about inpainting in the time-frequency domain, [^1] In other fields, one can go even further and require inpainting purely in the frequency domain, see for example [14], [15]. besides the fact that it is an interesting problem per se. In the case under consideration, a portion of time-frequency (TF) coefficients222 typically in the form of a complex-valued matrix is missing, which need to be recovered to allow a later reconstruction of the time-domain signal. Indeed, there have been a few methods developed in the TF context, all of them relying on neural modeling [16, 17, 18]. The most recent work [18] is based on the concept of a “deep prior”, a neural network that does not need training data besides the corrupted observation [19].
@@
II Gaps and the Short-time Fourier Transform
Note that a compact gap in the time domain is far from being equivalent to a gap in the TF domain, which can be associated with the uncertainty principle. When a time-domain signal is transformed using the Short-time Fourier Transform (STFT), it is first divided into overlapping frames, which are weighted using windows such as the Hann window (see, e.g., [20]). When a gap is present, the number of TF coefficients affected by the gap depends on how much the frames overlap. In practical scenarios, the gap overlaps with more than one frame; setting the missing samples to zeros then reduces the value of the corresponding TF coefficients, leading to a kind of TF-fadeout. For wide gaps, even a number of completely zero columns can appear in the TF domain (i.e., in the spectrogram), see the analysis in [21]. Since the inverse STFT also uses windowed frames, the same holds in the opposite direction: consecutive empty spectrogram columns result in a decrease in the time-domain amplitude, governed by the window shape. In turn, there is no way how to equitably compare time-domain inpainting techniques with those specialized for the TF-domain.
III Deep Prior Audio Inpainting (DPAI)
The deep prior, introduced in [19], is a concept used for the reconstruction of corrupted signals using a convolutional neural network (CNN). In contrast to most CNN applications, [19] does not utilize a training dataset; the reconstruction only involves a proper CNN architecture and the corrupted signal. The deep prior concept has also proven that the network architecture matters significantly more than previously thought.
III-A Application of Deep Prior to Audio Inpainting
@@
An encoder–decoder CNN denoted $f_{\theta}(\mathbf{n})$ is assumed to be able to generate any signal of size $N$ using a random but fixed input noise $\mathbf{n}\in\mathbb{R}^{N}$, using a proper $\theta$. This is a reliable assumption since the number of network parameters ${\theta}$ (weights and biases) is significantly greater than $N$.
In the deep prior reconstruction, also the parameters $\theta$ are initialized randomly. To obtain a reconstruction result, $\theta$ is then optimized as follows:
@@
where $E$ is a differentiable data fidelity function measuring the proximity of its two inputs. The optimal $\theta^{*}$ can be obtained using an iterative optimizer, e.g., Adam [22]. The crucial observation is that the optimization must be interrupted before it reaches $\theta^{*}$, for which we would have $f_{\theta^{*}\!}(\mathbf{n})=\mathbf{x}^{\textup{cor}}$, meaning basically overfitting. Instead, early stopping of the training is desired, resulting in an estimate $\hat{\theta}$ and a corresponding $\hat{\mathbf{x}}=f_{\hat{\theta}}(\mathbf{n})$ [19].
@@
The authors of [18] used an analog approach to inpaint audio in the time-frequency domain. Through the use of the STFT, denoted $\mathbf{L}$ in the following, an audio recording $\mathbf{x}$ is considered merely in the TF domain, where it forms a spectrogram matrix $\mathbf{L}\mathbf{x}=\mathbf{X}\in\mathbb{C}^{T\times F}$. Therefore the CNN operates on spectrograms and $\hat{\mathbf{X}}=f_{\hat{\theta}}(\mathbf{N})$ for a random but fixed matrix $\mathbf{N}$. Once $\hat{\mathbf{X}}$ has been established, the time-domain solution is obtained by the inverse STFT.
@@
Regarding the mask, in [18] only the case of completely missing columns of the spectrogram is considered, i.e., for the TF mask $\mathbf{M}\in\{0,1\}^{T\times F}$, each of its columns is either full of ones or zeros. Nevertheless, a generalization to other TF masks is straightforward. The observed spectrogram $\mathbf{X}^{\textup{cor}}$ is finally modeled as $\mathbf{X}^{\textup{cor}}\!=\mathbf{M}\odot\mathbf{X}^{\textup{orig}}$, where $\mathbf{X}^{\textup{orig}}$ corresponds to $\mathbf{x}^{\textup{orig}}$ in the time domain, i.e., $\mathbf{X}^{\textup{orig}}=\mathbf{L}\mathbf{x}^{\textup{orig}}$. Note that by analogy with the time-domain case, we obtain the feasible set $\mathit{\Gamma}_{\textup{TF}}=\{\mathbf{X}\,|\,\mathbf{M}\odot\mathbf{X}=\mathbf{M}\odot\mathbf{X}^{\textup{orig}}\}$.
In [18], the loss function $E$ in (1) linearly combines the classic Mean Squared Error between spectrograms and the multi-scale spectrogram loss [23, 24].
III-B Architecture
@@
The authors of [18] used an architecture similar to U-Net called MultiResUNet [25]. It is a common CNN architecture with several layers; the skip connections between the encoder and decoder stages play a significant role in audio inpainting. The architecture additionally includes the so-called harmonic convolution [26], which proved to be the best among multiple tested architectures in [18]. The harmonic convolution is a variation on the common 2D convolution that takes into account higher harmonics in audio.
III-C Variants
@@
The DPAI intrinsically reconstructs the whole $\hat{\mathbf{X}}$. Yet, considering that only a part of the spectrogram was not observed, the authors of [18] evaluate the performance of their method after injecting the reliable coefficients back to the solution as a postprocessing step. Formally, such an action is a projection
@@
onto the set $\mathit{\Gamma}_{\textup{TF}}$. Note that a resulting sharp transition between the regions is smeared in the time domain due to the nature of the STFT [21]; see also Sec. II. We call the reconstructions modified by (2) with context, while the others using the entire $\hat{\mathbf{X}}$ are without context. In the experimental part, we will study whether employing the context brings an improvement in performance.
IV Janssen spectrogram inpainting (Janssen-TF)
@@
$\sum_{i=1}^{p+1}a_{i}x_{n+1-i}=e_{n},\quad n=1,\dots,N+p.$ (3)
Here, $\mathbf{a}=[1,a_{2},\dots,a_{p+1}]^{\top}$ is the vector of model parameters (AR coefficients), $p$ is the model order, and $\mathbf{x}$ is zero-padded such that its final length is $N+p$. The error term $\mathbf{e}\in\mathbb{R}^{N+p}$ represents a realization of a zero-mean white noise process with variance $\sigma^{2}$ [27, Def. 3.1.2].
@@
Step 1 can be solved efficiently using standard methods, such as the Durbin–Levinson recursion (see for example the lpc function in Matlab). Step 2 boils down to a mean squares problem which can be treated e.g. via matrix inversion.
The proposed adaptation of the Janssen iteration to the TF setting, denoted Janssen-TF, lies in modifying step 2. In our case, the signal spectrogram is constrained to fit the observation, $\mathbf{X}^{\textup{cor}}$, at the reliable positions, while the signal itself follows the estimated AR model. To formalize this modification, note that the error in (3) can be written as $\mathbf{e}=\mathbf{A}\mathbf{x}$, where $\mathbf{A}\in\mathbb{R}^{(N+p)\times N}$ is a Toeplitz matrix composed using the (fixed) coefficients $\mathbf{a}$. We can then pose the proposed modification of step 2 as
@@
where $\mathbf{L}$ is the STFT and $\iota_{\mathit{\Gamma}_{\textup{TF}}}$ denotes the indicator function of the convex set $\mathit{\Gamma}_{\textup{TF}}$ of feasible spectrograms. [^3] The original Janssen algorithm minimizes the function $\frac{1}{2}\|\mathbf{A}\mathbf{x}\|+\iota_{\mathit{\Gamma}_{\textup{T}}}(\mathbf{x})$ in subproblem 2. As such, the problem (4) is convex and it is solvable using standard numerical methods. We propose to employ the alternating direction method of multipliers (ADMM), which includes two principal steps [28, Sec. 3.1.1]:
@@
The whole algorithm is summarized in Alg. 1. For the simplification of (5a), we have posed the assumption [^4] This assumption is met for instance if $\mathbf{L}$ corresponds to a Parseval tight frame, which can be ensured by a suitable practical setting of the STFT. $\mathbf{L}^{*}\mathbf{L}=\mathbf{I}$, leading to $\bar{\mathbf{x}}^{(k+1)}=\operatorname{prox}_{\frac{1}{2}\|\cdot\|^{2}}(\mathbf{L}^{*}(\mathbf{z}^{(k)}-\mathbf{u}^{(k)}))$, see [29, Rem. 3], which then corresponds to line 1 of Alg. 1 [30, Table 1]. Similarly, (5b) boils down to line 1 of the algorithm; this is computed via (2). The hyperparameters of the algorithm are the number of outer and inner iterations ($I$ and $K$, respectively), and the step size $\rho>0$.
@@ 11      $\mathbf{x}^{(i+1)}=\bar{\mathbf{x}}^{(K+1)}$
Note that several other algorithms can be used to solve (4) numerically, such as the primal–dual algorithm [31] or the Douglas–Rachford or the forward–backward algorithms [30], combined with the approximal operator of the function $\iota_{\mathit{\Gamma}_{\textup{TF}}}(\mathbf{L}\,\cdot)$ [32]. However, the variant with ADMM described above, denoted Janssen-TF-ADMM, exhibited the best performance out of six algorithms in our preliminary testing.
V Experiments
V-A Data, setup and metrics
@@
The STFT (i.e., the operator $\mathbf{L}$) utilizes a 2048-sample-long Hann window (corresponding to 128 ms [^6] This exceeds the standard window length for speech processing; however, a change of this hyperparameter would cause the need for changing the DPAI architecture.), with 75% overlap and 2048 frequency channels; this setting is the default for the time-frequency transform of the librosa package [^7] https://librosa.org/doc/0.10.2/generated/librosa.stft.html, as used in [18], and ensures that $\mathbf{L}$ forms a Parseval tight frame.
@@
  • gap-wise Janssen method in the time domain [12].
The gap-wise method is a recent twist to the original Janssen algorithm [2], which is shown in [12] to be the state-of-the-art time-domain inpainter. As explained in Sec. II, the time-domain Janssen method is not directly comparable with the first two. We use it such that we transform the corrupted spectrogram $\mathbf{X}^{\textup{cor}}$ to the time-domain and zero-out all samples affected by the zero columns of $\mathbf{X}^{\textup{cor}}$.
@@
V-B Results
The objective comparison on the DPAI dataset is shown in Fig. 1. The first observation is that the time-domain-based Janssen method [12] is not competitive with the other approaches. While the average performance of the DPAI variants mostly lies within the confidence intervals of each other, suggesting limited significance of the difference, the proposed Janssen-TF-ADMM method largely outperforms the competitors in both SNR and ODG (except the longest gaps).
11112222333344445555666600101010102020202030303030404040405050505060606060gap lengthSNR (dB)DPAI without contextDPAI with contextgap-wise JanssenJanssen-TF-ADMM
1111222233334444555566663.53.5-3.5- 3.533-3- 32.52.5-2.5- 2.522-2- 21.51.5-1.5- 1.511-1- 10.50.5-0.5- 0.500gap lengthODGDPAI without contextDPAI with contextgap-wise JanssenJanssen-TF-ADMM
Figure 1: Comparison of the inpainting methods using objective metrics – SNR (top) and ODG (bottom). The plots show results averaged over the 8 test signals together with the bootstrap interval estimates of the mean values at the $\alpha=5\%$ significance level [37, p. 160]. If the intervals do not overlap, it may be concluded that the difference of the means is statistically significant.
To support the objective evaluation on the first dataset, Fig. 2 presents the results of the subjective listening test. Its results correlate well with the objective evaluation; in particular, Janssen-TF-ADMM is superior to the other methods. Moreover, this result is statistically significant in terms of medians. The DPAI variants scored almost the same.
anchorreferenceDPAI without contextDPAI with contextgap-wise JanssenJanssen--TF-ADMM0020202020404040406060606080808080100100100100MUSHRA score